Circular No. 5746

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 T. J. Balonek and C. A. Tremonti, Colgate University; and J. Koch,

St. Lawrence University, report: “V -band CCD photometry with the Burrell Schmidt telescope, Kitt Peak, indicates a near-linear daily decrease in brightness of 0.30 mag during Mar. 31.2-Apr. 2.4 UT. Using GSC 4383.0928 and GSC V = 11.42, we obtained Mar. 31.17 UT, 10.31; 31.26, 10.34; Apr. 1.21, 10.59; 1.40, 10.65; 2.13, 10.89; 2.22, 10.91; 2.37, 10.97. Each value is the average of three to five consecutive exposures. A different detector was used for the Apr. 1 observations; color corrections have not yet been applied to these measurements.”

It should be noted that Corwin’s reference-star V magnitudes on IAUC 5742 round to 0.1 mag to those in the Thompson-Bryan Supernova Search Charts and Handbook, which also gives V = 14.0 for GSC 4383.0340.

HDE 245770 C. Bartolini, G. Cosentino, A. Guarnieri and A. Piccioni, Bologna Uni-

versity, communicate: “Spectroscopic observations of the Be x-ray transient HDE 245770 = A0535+26 with the Bologna Observatory’s 1.5-m telescope on Mar. 31.8 UT and Apr. 1.8 reveal enhanced activity. Compared to the situation on Mar. 9.8 Hα and Hβ are strongly in emission. Hγ, which is usually seen in absorption, is also in emission, and so are He I 587.6-nm and 667.8-nm. X-ray monitoring of the system is encouraged.”

V1974 CYGNI (NOVA CYGNI 1992) P. Garnavich, Dominion Astrophysical Observatory, reports: “Spectra

obtained at the DAO between Mar. 26 and 29 clearly show the presence of [Fe X] emission at 637.4 nm. The [Fe X] flux is 5 percent of the Hα strength. Coronal lines of [Ar X] 553.2 nm and possibly of [Fe XIV] 530.4 nm (blended with O VI 529.1 nm) are also seen. Such high ionization potentials are not normally observed in classical novae. Nebular lines of [O I] 630.0 and [O I] 575.5 nm are also present, and their profiles consist of many sharp components unresolved at our 0.17-nm resolution. All the lines are broad with measured FWHM of 2000 km/s. The range of observed ionization states is similar to that seen in the recurrent nova RS Oph (Wallerstein and Garnavich 1986, PASP 98, 875) and suggests that the coronal lines are a result of strong shocks. Observations at x-ray, extreme-ultraviolet and ultraviolet wavelengths are suggested.”

1993 April 4 Brian G. Marsden